# The Language of Interoception: Examining Embodiment and Emotion Through a Corpus of Body Part Mentions

### v1

@@ Appendix G Controling for Post Length (Supplementary Tables, BA2)
Appendix H Emotion associations between specific “my <BPM>" types (Supplementary Figures, BA3)
@@
In this section, we include exact values for the differences in the percentage of sentences with emotion-associated words in samples containing "my <BPM>" types in. These are shown in Figures 13 (TUSCctry-BPM and VAD), 14 (TUSCctry®-BPM and emotion categories), 16 (Spinn3rBPM and VAD), and 15 (Spinn3rBPM and emotion categories). We focus on the top 30 most common “my <BPM>" types in TUSCcity-BPM. and the and top 15 most common body parts in Spinn3rBPM. The mean and standard deviation is also calculated over all common “my <BPM>" types analyzed for each corpus for each emotional dimension, and for each “my <BPM>" type we display the ’delta’ as the proportion of the “my <BPM>" type sample co-occuring with an emotion-associated words subtracted from the mean. All word-emotion associations are from the NRC VAD lexicon and the NRC emotion lexicon.
@@
Figure 13: BA2 - TUSCctry-myBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>" types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>" mean) in the percentage of tweets with at least one word that is associated with high/low valence, arousal, and dominance (according to the NRC VAD lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>" types present in the dataset).
@@
Figure 14: BA2 - TUSCctry-myBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>" types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>" mean) in the percentage of tweets with at least one word that is associated with each emotional category (according to the NRC emotion lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>" types present in the dataset).
@@
Figure 15: BA2 - Spinn3rmyBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>" types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>" mean) in the percentage of blog sentences with at least one word that is associated with each emotional category (according to the NRC emotion lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>" types present in the dataset).
@@
Figure 16: BA2 - Spinn3rmyBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>" types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>" mean) in the percentage of blog sentences with at least one word that is associated with high/low valence, arousal, and dominance (according to the NRC VAD lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>" types present in the dataset).
@@
Appendix I How are body part words represented in word-emotion association lexicons? (Additional Experiment)
@@
Category High V. (%) Low V. (%) High A. (%) Low A. (%) High D. (%) Low D. (%)
NoBPM 8.38 1.69 1.89 4.22 3.59 1.92
BPM 8.12 2.09 2.02 4.71 3.34 2.36
YourBPM 8.85 2.20 2.16 4.42 3.49 2.27
MyBPM 7.98 2.61 2.11 4.68 2.97 2.78
3pBPM 7.41 2.58 2.21 4.36 3.23 2.78
Table 6: BA2 - Spinn3r - Average percentage of words in each sample associated with high/low VAD based on NRC VAD lexicon for each BPM category. For each sample, the proportion is calculated as the number of words associated in a given category divided by the total number of words in the sample that appear in the lexicon.
Category Anger (%) Fear (%) Joy (%) Sadness (%) Surprise (%) Trust (%)
No BPM 0.66 0.77 1.35 0.76 0.67 1.50
<BPM> 0.81 0.98 1.32 0.97 0.70 1.48
your <BPM> 0.87 1.07 1.57 0.91 0.83 1.51
my <BPM> 0.95 1.19 1.41 1.30 0.77 1.30
her/his/their <BPM> 1.04 1.28 1.33 1.28 0.71 1.61
Table 7: BA2 - Spinn3r - Average percentage of words in each sample associated with emotional categories based on NRC emotion lexicon for each BPM category. For each sample, the proportion is calculated as the number of words associated in a given category divided by the total number of words in the sample that appear in the lexicon.
Appendix J How does referring to one’s own body change the emotional signature in personal narratives online (BA1: Supplementary Figures)
Table 8 displays the percentage of BPM vs no BPM sentences across High/Low VAD categories containing at least one word associated with each emotional category according to NRC VAD lexicon in the Spinn3r dataset across different bins of sentence length for high/low VAD, and Table 9 displays this data for emotional categories.
@@
Appendix K What are the emotions most commonly associated with the most frequently discussed body parts?
In Table 10, we display the top emotion (with associated increase in emotion-associated word from “my <BPM>" average) for top “my <BPM>" types across Spinn3rBPM and TUSCBPM datasets. A dash indicates the body part is not present in top “my <BPM>" types considered for the dataset (top 15 for Spinn3rBPM and top 30 for TUSCBPM).
@@
Table 10: Most associated emotion (with associated increase in emotion-associated word from “my <BPM>" average) for “my <BPM>" types across Spinn3rBPM and TUSCBPM datasets (top 15 for Spinn3rBPM and top 30 for TUSCBPM). A dash indicates the body part is not present in top “my <BPM>" types considered for the dataset.
Appendix L Is physical wellbeing correlated with emotional word use? (Additional Experiment)
We also look at whether the physical wellbeing indicators we examine in other experiments are correlated with emotion-related words according to the NRC lexicon. See Table 11.
@@
Appendix M Why/when do we refer to our own bodies? (Additional Experiment)
Method: We evaluate the context of words that tend to surround myBPMs by looking at word clouds which visualize the words which are most likely to appear within the context window of particular myBPMs (See Figure 17).
Results: We find that there are significantly more 3pBPM types with $>0.1\%$ occurrence compared to myBPM types with $>0.1\%$ (131 vs 57 in the Spinn3rBPM dataset, and 108 vs 56 in the TUSCctry-BPM dataset). We also find that, although myBPM types exhibit a rich diversity in associated contexts, that some myBPM types share common contexts as well, especially "hurt", "pain", and "sick", which frequently co-occur with several frequent myBPMs such as "my head", "my back", "my neck", and" my stomach".
Discussion: The analysis reveals that third-person BPM types (3pBPM) in are significantly more diverse than myBPM types at the 0.1% occurrence threshold in the Spinn3r dataset, indicating a more limited and concentrated vocabulary when people refer to their own body than the bodies of others. The words with negative associations with health frequently accompanying some of the most common myBPM types also highlight health concerns and physical pain as central themes for myBPM usage.
Figure 17: B6 - Wordclouds for the twenty most frequent myBPM types in the TUSCcity dataset with the most frequent co-occurring words.

### v2

@@ Appendix H Effect of time on BPM usage (Supplementary statistical tests, B4)
Appendix I Emotion associations between specific “my <BPM>” types (Supplementary Figures and Statistical Tests, BA3)
@@
In this section, we include exact values for the differences in the percentage of sentences with emotion-associated words in samples containing "my <BPM>" types in. These are shown in Figures 13 (TUSCctry-BPM and VAD), 14 (TUSCctry®-BPM and emotion categories), 15 (Spinn3rBPM and emotion categories), and 16 (Spinn3rBPM and VAD). We focus on the top 30 most common “my <BPM>” types in TUSCcity-BPM. and the and top 15 most common body parts in Spinn3rBPM. The mean and standard deviation is also calculated over all common “my <BPM>” types analyzed for each corpus for each emotional dimension, and for each “my <BPM>” type we display the ’delta’ as the proportion of the “my <BPM>” type sample co-occuring with an emotion-associated words subtracted from the mean. All word-emotion associations are from the NRC VAD Lexicon and the NRC Emotion Lexicon.
We also use a two-way ANOVA test to examine the independent and interactive effects of BPM category and emotional category on VAD (Valence-Arousal-Dominance) values, which revealed significant main effects for both category (F(3,48) = 26.94, p < 0.001) and emotion (F(5,48) = 60.72, p < 0.001), with emotion showing the stronger effect, but no significant interaction between factors (F(15,48) = 1.25, p = 0.267). This indicates that while values vary across the different emotions (as expected, since some emotion-associated words naturally occur in certain corpora more frequently), the different BPM instance categories independently have a statistically significant effect emotion-associated word co-occurences. We also evaluate this test for the different emotional categories, again showing significant effects of the BPM categories (F(3,64) = 54.33, p < 0.001) and emotion (F(7,64) = 44.68, p < 0.001), and in this case also a significant interaction between category and emotion (F(21,64) = 2.03, p = 0.016) – suggesting that the impact of BPM category is statistically significant, and also varies in strength depending on the specific emotion dimension.
@@
Figure 13: BA2 - TUSCctry-myBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>” types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>” mean) in the percentage of tweets with at least one word that is associated with high/low valence, arousal, and dominance (according to the NRC VAD lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>” types present in the dataset).
@@
Figure 14: BA2 - TUSCctry-myBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>” types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>” mean) in the percentage of tweets with at least one word that is associated with each emotional category (according to the NRC Emotion Lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>” types present in the dataset).
@@
Figure 15: BA2 - Spinn3rmyBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>” types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>” mean) in the percentage of blog sentences with at least one word that is associated with each emotional category (according to the NRC Emotion Lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>” types present in the dataset).
@@
Figure 16: BA2 - Spinn3rmyBPM - Variance in emotion-associated term co-occurrence for top 30 most common “my <BPM>” types present in the dataset. For each type, we display the delta in ("my <BPM>" type minus “my <BPM>” mean) in the percentage of blog sentences with at least one word that is associated with high/low valence, arousal, and dominance (according to the NRC VAD lexicon). Mean and standard deviation are calculated over all body parts considered (top 30 most common “my <BPM>” types present in the dataset).
@@
Appendix J How are body part words represented in word-emotion association lexicons? (Additional Experiment)
@@
Appendix K How does referring to one’s own body change the emotional signature in personal narratives online (BA1: Supplementary Figures)
Table˜6 displays the percentage of BPM vs no BPM sentences across High/Low VAD categories containing at least one word associated with each emotional category according to NRC VAD lexicon in the Spinn3r dataset across different bins of sentence length for high/low VAD, and Table˜7 displays this data for emotional categories.
@@
Appendix L What are the emotions most commonly associated with the most frequently discussed body parts?
In Table˜8, we display the top emotion (with associated increase in emotion-associated word from “my <BPM>” average) for top “my <BPM>” types across Spinn3rBPM and TUSCBPM datasets. A dash indicates the body part is not present in top “my <BPM>” types considered for the dataset (top 15 for Spinn3rBPM and top 30 for TUSCBPM).
@@
Table 8: Most associated emotion (with associated increase in emotion-associated word from “my <BPM>” average) for “my <BPM>” types across Spinn3rBPM and TUSCBPM datasets (top 15 for Spinn3rBPM and top 30 for TUSCBPM). A dash indicates the body part is not present in top “my <BPM>” types considered for the dataset.
Appendix M Is physical wellbeing correlated with emotional word use? (Additional Experiment)
We also look at whether the physical wellbeing indicators we examine in other experiments are correlated with emotion-related words according to the NRC lexicon. See Table˜9.
@@
Appendix N Why/when do we refer to our own bodies? (Additional Experiment)
Method: We evaluate the context of words that tend to surround myBPMs by looking at word clouds which visualize the words which are most likely to appear within the context window of particular myBPMs (See Figure˜17).
Results: We find that there are significantly more 3pBPM types with $0.1\%$ occurrence compared to “my <BPM>” types with $0.1\%$ (131 vs 57 in the Spinn3rBPM dataset, and 108 vs 56 in the TUSCctry-BPM dataset). We also find that, although “my <BPM>” types exhibit a rich diversity in associated contexts, that some “my <BPM>” types share common contexts as well, especially "hurt", "pain", and "sick", which frequently co-occur with several frequent myBPMs such as "my head", "my back", "my neck", and" my stomach".
Discussion: The analysis reveals that third-person BPM types (3pBPM) in are significantly more diverse than “my <BPM>” types at the 0.1% occurrence threshold in the Spinn3r dataset, indicating a more limited and concentrated vocabulary when people refer to their own body than the bodies of others. The words with negative associations with health frequently accompanying some of the most common “my <BPM>” types also highlight health concerns and physical pain as central themes for myBPM usage.
Figure 17: B6 - Wordclouds for the twenty most frequent “my <BPM>” types in the TUSCcity dataset with the most frequent co-occurring words.